OpenStudy (idealist10):

Solve y'+y=(2xe^-x)/(1+ye^x) using variation of parameters followed by separation of variables.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

dy/dx=(2xe^-x)/(1+ye^x)-y?

OpenStudy (anonymous):

So you have to solve the same equation twice, using a different method each time?

OpenStudy (idealist10):

Same equation twice? I don't think so...

OpenStudy (anonymous):

I guess I just don't understand what the directions mean by "...followed by..." Anyway, to use VoP, you'll need two known fundamental solutions. From the homogeneous equation \[y'+y=0\] you have the characteristic equation \[r+1=0~~\iff~~r=-1\] So one solution is \(C_1e^{-x}\) (the fundamental solution is then \(e^{-x}\)). Not sure about the second solution just yet...

OpenStudy (idealist10):

I'll come back. Let me look for problems like this.

OpenStudy (anonymous):

I don't know if it's relevant or helpful, but I stumbled across an interesting way to express the right hand side in terms of partial fractions: \[\begin{align*}\frac{xe^{-x}}{1+ye^x}&=\frac{x}{e^x(1+ye^x)}\\\\ &=\frac{f(x,y)}{e^x}+\frac{g(x,y)}{1+ye^x}\\\\ x&=f(x,y)(1+ye^x)+g(x,y)e^x\\\\ x&=f(x,y)+e^x\bigg(yf(x,y)+g(x,y)\bigg) \end{align*}\] which means \(f(x,y)=x\) and \(yf(x,y)+g(x,y)=0\). Substituting gives \(g(x,y)=-xy\), and so \[RHS=\frac{xe^{-x}}{1+ye^x}=\frac{x}{e^x}-\frac{xy}{1+ye^x}\]

OpenStudy (anonymous):

So now you can split up the equation into two non-homogeneous ones: \[y'+y=xe^{-x}~~~~~~~~~~~\text{(a simple linear equation)}\\ y'+y=-\frac{xy}{1+ye^x}~~\text{(not quite sure yet)}\]

OpenStudy (anonymous):

Mathematica 9 came up with the follwoing:\[y(x)=-e^{-x}\pm e^{-2 x} \sqrt{c_1 e^{2 x}+2 e^{2 x} x^2+e^{2 x}} \]

OpenStudy (anonymous):

The linear equation gives \(e^{-x}x^2\) as a fundamental solution (I'll leave the details of finding it to you). This solution shows up under the square root in rob's comment

OpenStudy (idealist10):

Thanks for the extra info.

OpenStudy (anonymous):

You're welcome! So when you use VoP, you'll be using \(y_1=e^{-x}\) and \(y_2=x^2e^{-x}\), which should be linearly independent... Check to make sure that's true.

OpenStudy (anonymous):

I'm not sure where the separation of variables comes in though...